# Identifying Weight-Variant Latent Causal Models

### v6

@@ 3.2 Transitivity: the Challenge of Identifying Causal Representations
Even with the identifiable $n_{i}$, it is impossible to identify the latent causal variables $z_{i}$ without additional assumptions. To interpret this point, for simplicity, let us only consider the influences of $z_{1}$ and $z_{2}$ on $\mathbf{x}$ in Figure 1. According to the graph structure in Figure 1, assume that $z_{1}:=n_{1}$, $z_{2}:=z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})+{\boldsymbol{\varepsilon}}$ (case 1). We then consider the graph structure shown in the left column of Figure 2, where the edge $z_{1}\rightarrow z_{2}$ has been removed, and assume that $z_{1}:=n_{1}$, $z_{2}:=n_{2}$ and $\mathbf{x}:=\mathbf{f}\circ\mathbf{g}(z_{1},z_{2})+{\boldsymbol{\varepsilon}}$ where $\mathbf{g}(z_{1},z_{2})=[z_{1},z_{1}+z_{2}]$ (case 2). Interestingly, we find that the causal models in case 1 and case 2 generate the same observed data $\mathbf{x}$, which implies that there are two different causal models to interpret the same observed data. Clearly, $z_{2}$ in both the two equivalent causal structures are different, and thus $z_{2}$ is unidentifiable.
@@ 3.3 Scaling Indeterminacy in Causal Representations
The scaling indeterminacy of latent causal variables is also an intrinsic indeterminacy in latent space. Again, for simplicity, we only consider the influence of $z_{1}$ and $z_{2}$ on $\mathbf{x}$ in Figure 1, and assume that $z_{1}:=n_{1}$, $z_{2}:=z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})$. Under this setting, if the value of $z_{1}$ is scaled by $s$, e.g. $s\times z_{1}$, we can easily obtain the same observed data $\mathbf{x}$ by: 1) letting $z_{2}:=\frac{1}{s}\times z_{1}+n_{2}$ and 2) $\mathbf{x}:=\mathbf{f}\circ\mathbf{g}(z_{1},z_{2})$ where $\mathbf{g}(z_{1},z_{2})=[s\times z_{1},z_{1}]$. This indeterminacy is because the scaling of the latent variables $z_{i}$ can be ‘absorbed’ by the nonlinear function from $\mathbf{z}$ to $\mathbf{x}$ and the causal functions among the latent causal variables. Therefore, without additional information to determine the values of the latent causal variables $z_{i}$, it is only possible to identify the latent causal variable up to scaling, not exactly recovering the values. In general, this scaling does not affect identifying the causal structure among the latent causal variables. We will further discuss this point in Section 4.
@@ 3.4 Permutation Indeterminacy in Causal Representations
However, the flexibility and ambiguity above in latent space also offer an advantage, i.e., by pre-defining a causal order, we can avoid constraints like those proposed in Zheng et al. (2018), facilitating the learning of latent causal representations and causal structure. For example, For example, assume that the true latent causal relations for generating data are: size $\rightarrow$ color $\rightarrow$ shape $\rightarrow$ location. During the learning process, we can pre-define a causal order $z_{1},z_{2},z_{3},z_{4}$, without specifying semantic information for the nodes. Since $z_{1}$ is the first node in the predefined causal order, $z_{1}$ is enforced to the semantic information of the first node in the true underlying causal order, e.g., the size. Similarly, $z_{2}$ is enforced to the semantic information of the second node in the true underlying causal order, e.g., the color. Furthermore, the predefined causal order allows us to naturally establish a causal fully-connected graph as depicted in Figure 3, ensuring the estimation of a directed acyclic graph (DAG) in learning causal representation and avoiding DAG constraints, such as those proposed by Zheng et al. (2018). We will further demonstrate how to implement this in the proposed method in Section 6.
@@ 4 Identifiable Causal Representations with Weight-variant Linear Gaussian Models
As discussed in Section 3.2, the key factor that impedes identifiable causal representations is the transitivity in latent space. Note that the transitivity is because the causal influences among the latent causal variables may be ‘absorbed’ by the nonlinear mapping from latent variables $\mathbf{z}$ to the observed variable $\mathbf{x}$. To address this issue, motivated by identifying causal structures with the change of causal influences in observed space (Ghassami et al., 2018; Huang et al., 2019, 2020), we allow causal influences among latent causal variables to be modulated by an additionally observed variable $\mathbf{u}$, represented by the red edge in Figure 4.
Figure 4: The proposed Latent Causal Model. Here we use the ’red’ lines to indicate the changes of weights across $\mathbf{u}$, which is different from the definition of edges (e.g., causal direction) in standard SCM.
Intuitively, variant causal influences among latent causal variables cannot be ‘absorbed’ by an invariant nonlinear mapping from $\mathbf{z}$ to $\mathbf{x}$, breaking the transitivity, resulting in identifiable causal representations. Specifically, we explore latent causal generative models where the observed data $\mathbf{x}$ is generated by the latent causal variables $\mathbf{z}$, allowing for any potential graph structures among $\mathbf{z}$. In addition, there exist latent noise variables $\mathbf{n}$, known as exogenous variables in causal systems, corresponding to latent causal variables. We introduce a surrogate variable $\mathbf{u}$ characterizing the changes in the distribution of $\mathbf{n}$, as well as the causal influences among latent causal variables $\mathbf{z}$. Here $\mathbf{u}$ could be environment, domain, or time index. More specifically, we parameterize the causal generative models as follows:
@@
  • In Eq. (2), $\boldsymbol{\lambda_{i}}(\mathbf{u})$ denote the vector corresponding to the causal weights from $\mathbf{z}$, $\mathbf{z}\in\mathbb{R}^{n}$, to $z_{i}$, e.g. $\boldsymbol{\lambda_{i}}(\mathbf{u})=[\lambda_{1,i}(\mathbf{u}),\lambda_{2,i}(\mathbf{u}),...,\lambda_{n,i}(\mathbf{u})]$, and each ${\lambda_{j,i}}$ could be a nonlinear mapping.
@@
In Eq. (2), we assume the latent causal model of each $z_{i}$ satisfies a linear causal model with the causal weights being modulated by $\mathbf{u}$, and we call it weight-variant linear Gaussian model. Therefore, $p({\mathbf{z}|\mathbf{u}})$ satisfies the following multivariate Gaussian distribution:
@@
where $\mathrm{pa}_{i}$ denotes the parent nodes of $z_{i}$. Here, $\mathbf{u}$ can represent the domain index, condition index, or time index. Different values of $\mathbf{u}$ give different causal models of $\mathbf{z}$, resulting in nonstationary data of $\mathbf{z}$, so the modulation of $\mathbf{u}$ can be considered as soft interventions done by nature. Making use of distribution shifts to identify causal structures has been widely considered for causal structure learning among observed variables (Ghassami et al., 2018; Huang et al., 2019, 2020; Schölkopf et al., 2021), and here we extend it to identify latent causal structures.
Furthermore, this multi-variate Gaussian can be re-formulated with the following exponential family:
@@
where the parameter $\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u})=[{\Sigma}^{-1}\boldsymbol{\mu};-\frac{1}{2}\operatorname{vec}({{\Sigma}^{-1}})]$, $Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u})$ denotes the normalizing constant, and the sufficient statistic $\mathbf{T}_{\mathbf{z}}(\mathbf{z})=[\mathbf{z};\operatorname{vec}(\mathbf{z}\mathbf{z}^{T})]$ (‘vec’ denotes the vectorization of a matrix). We further denote by $\mathbf{T}_{\text{min}}$ the minimal sufficient statistic and by $k$ its dimension, with $22n\leq k\leq n+(n(n+1))/$ 2. In particular, $k=2n$ corresponds to the case when $\mathbf{T}_{\text{min}}(\mathbf{z})=[{z_{1}},z_{2},...,z_{n},{z^{2}_{1}},{z^{2}_{2}},...,{z^{2}_{n}}]$, or in other words, there is no edges among the latent variables $\mathbf{z}$, while $k=n+(n(n+1))/2$ corresponds to a full-connected causal graph over $\mathbf{z}$. So with different causal structures over $\mathbf{z}$, the dimension $k$ may vary. For the graph structure in Figure 1, $\mathbf{T}_{\text{min}}(\mathbf{z})=[{z_{1}},z_{2},z_{3},z_{4},{z^{2}_{1}},{z^{2}_{2}},{z^{2}_{3}},{z^{2}_{4}},z_{1}z_{2},z_{1}z_{3},z_{3}z_{4}]$ and $k=11$.
The following theorem shows that under certain assumptions on the nonlinear mapping $\mathbf{f}$ and the variability of $\mathbf{u}$, the latent variable $\mathbf{z}$ can be identifiable up to trivial permutation and scaling.
@@
    There exist $12n+$ 1 distinct points $\mathbf{u}_{\mathbf{n},0},\mathbf{u}_{\mathbf{n},1},...,\mathbf{u}_{\mathbf{n},2n}$ such that the matrix
@@
then the true latent causal variables $\mathbf{z}$, which are learned by matching the true marginal data distribution $p(\mathbf{x}|\mathbf{u})$, are related to the estimated latent causal variables ${\mathbf{\hat{z}}}$ by the following relationship: $\mathbf{z}=\mathbf{P}\mathbf{\hat{z}}+\mathbf{c},$ where $\mathbf{P}$ denotes the permutation matrix with scaling, $\mathbf{c}$ denotes a constant vector.
@@
The proof can be done according to the following intuition. With the support of assumptions (i)-(iii), one can utilize the identifiability result from nonlinear ICA to identify the latent noise variables $\mathbf{n}$ up to permutation and scaling, e.g., $\mathbf{n}=\mathbf{P}\mathbf{\hat{n}}+\mathbf{c}_{\mathbf{n}}$ where $\mathbf{\hat{n}}$ denotes the recovered latent noise variables obtained by matching the true marginal data distribution, and $\mathbf{c}_{\mathbf{n}}$ notes a constant vector. Next, using a similar technique, we can utilize assumption (iv) to establish a linear transformation between the true latent causal variables $\mathbf{z}$ and the recovered latent causal variables $\mathbf{\hat{z}}$, e.g., $\mathbf{z}=\mathbf{A}(\mathbf{\hat{z}})+\mathbf{c}_{\mathbf{z}}$ where $\mathbf{c}_{\mathbf{z}}$ denotes a constant vector, and $\mathbf{A}$ denotes a liner transformation. Finally, combining the identifiability result for the latent noise variables $\mathbf{n}$ with assumption (v), the linear transformation can be further reduced to permutation and scaling, e.g., $\mathbf{z}=\mathbf{P}(\mathbf{\hat{z}})+\mathbf{c}_{\mathbf{z}}$. Details can be found in Appendix A.1.
@@
Assumptions (i)-(iii) are originally developed by nonlinear ICA Hyvarinen and Morioka (2016); Hyvarinen et al. (2019); Khemakhem et al. (2020); Sorrenson et al. (2020). We here consider unitizes these assumptions considering the following two main reasons. 1) These assumptions have been verified to be practicable in diverse real-world application scenarios Kong et al. (2022); Xie et al. (2022b). 2) Our result eliminates the need to make assumption about the dimensionality of latent causal or noise variables, which is in contrast to existing methods that require prior knowledge of the dimensionality, due to imposing the two-parameter Gaussian noise variables Sorrenson et al. (2020). Assumptions (iv)-(v) originally introduced by this work. Intuitively, assumption (iv) is similar to assumption (iii) in the hope that the auxiliary variable $\mathbf{u}$ must have a sufficiently strong and diverse effect on the distributions of latent causal variables $\mathbf{z}$, similar to latent noise variables $\mathbf{n}$.
@@
Among the assumptions above, the only one that cannot be considered weak or natural is assumption (v), which was originally introduced in this work. Essentially, it is to ensure sufficient changes in causal influence. For simplicity, consider two variables, $z_{1}$ and $z_{2}$, and $z_{2}:=\lambda(\mathbf{u})z_{1}+n_{2}$, where $\lambda(\mathbf{u})=\hat{\lambda}(\mathbf{u})+b$, with $\lambda$ and $\hat{\lambda}$ belonging to the same function class. As a consequence, while the causal influence $\lambda(\mathbf{u})$ changes as a whole across $\mathbf{u}$, there always exists a part $bz_{1}$ that remains unchanged across $\mathbf{u}$. This unchanged term $bz_{1}$ across $\mathbf{u}$ can be absorbed into $\mathbf{f}$ the mapping from $\mathbf{z}$ to $\mathbf{x}$, resulting in a possible solution $\hat{z}_{2}:=\hat{\lambda}(\mathbf{u})z_{1}+n_{2}$, instead of the groundtruth $(\hat{\lambda}(\mathbf{u})+b)z_{1}+n_{2}$, which leads to an unidentifiable outcome. Essentially, this unidentifiable result occurs because a part of a weight remains unchanged across $\mathbf{u}$, a situation we refer to as insufficient change. The objective of assumption (v) is to prevent such insufficient changes. Specifically, by constraining the function class of $\lambda$, e.g., $\lambda(\mathbf{u}=0)=0$ and thus $\hat{\lambda}(\mathbf{u}=0)=0$, $\lambda(\mathbf{u})$ can not be expressed by $\hat{\lambda}(\mathbf{u})+b$ with $b\neq 0$. Intuitively, assumption (v) suggests that there is a scenario where an existing edge can be removed at a certain $\mathbf{u}$, e.g., $\lambda(\mathbf{u}=\mathbf{0})=0$. By removing the edge, we can ensure that no part of the weight corresponding to the edge remains unchanged. It is important to emphasize that assumption (v) is to limit the function class, so that $\lambda$ do not include a constant term, such as $b$, by assuming $\lambda(\mathbf{u}=0)=0$. Once samples are drawn from this limited function class, the assumption is met, making the observed data generated by these samples effective for inferring latent causal variables. Therefore, we do not need the samples to necessarily include the specific point $\mathbf{u}=0$, to generate the corresponding observed data, which is then used for inferring the latent causal variables.
Assumption (v) as a Bridge Between Soft and Hard Interventions for Identifiability
Interestingly, assumption (v) suggests that even if an edge exists between two latent causal variables, there can be a scenario where this edge is effectively removed for a specific $\mathbf{u}$, e.g., when $\lambda(\mathbf{u}=\mathbf{0})=0$. This concept closely aligns with the idea of hard interventions, where manipulating a variable can sever causal connections. This insight draws a compelling link between this work and recent advances that leverage hard interventions to improve the identifiability of causal representations (Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023). However, unlike these works, which typically rely on setting $\mathbf{u}=\mathbf{0}$, this work, through assumption (v), defines a restricted class of functions for $\lambda(\mathbf{u})$. By sampling from this restricted function class, identifiability can be achieved without necessarily requiring $\mathbf{u}=\mathbf{0}$, thereby broadening the applicability beyond existing hard intervention methods. Further exploration of this connection could pave the way for a more general framework that unifies soft and hard interventions for identifiability. However, such an investigation lies beyond the scope of this work.
@@
Assumption (v), for instance, could arise in the analysis of imaging data of cells, where various batches of cells are exposed to different small-molecule compounds (e.g., $\mathbf{u}$). In this scenario, each latent variable (e.g., $\mathbf{z}$) represents the concentration level of a distinct group of proteins, with protein-protein interactions (e.g., latent causal relations among $\mathbf{u}$) playing a significant role. There is considerable interest in computational methods aimed at extracting biological information from images (e.g., $\mathbf{x}$), given that imaging data offers a more cost-effective alternative to other assays while still yielding valuable insights into cell states Chandrasekaran et al. (2021). Research has revealed that the mechanisms of action of small molecules exhibit variations in selectivity Forbes and Krueger (2019), which can profoundly affect protein-protein interactions. In the context above, assumption (v) can be justifiable as certain small molecules have been demonstrated to significantly influence protein-protein interactions, thereby altering the dependency of the concentration of the products of these interactions on the concentration of the reactants. The assumption require that there exist $\lambda_{i,j}(\mathbf{u}=\mathbf{0})=0$. This implies that the causal relationship may be disconnected. In the context above, it is possible for a small molecule compound to disrupt or inhibit protein-protein interactions (PPIs), effectively causing the interaction to disappear Arkin and Wells (2004). Such small molecules are often referred to as inhibitors or disruptors of protein-protein interactions. These compounds can work through various mechanisms, such as blocking the binding interface between the two proteins, destabilizing the protein complex, altering the conformation of one or both proteins involved in the interaction, or interfering with the signaling pathways associated with the interaction. The development of small molecule inhibitors of PPIs is an active area of research in drug discovery, especially in diseases where dysregulated protein interactions contribute to pathology Lu et al. (2020); Bojadzic et al. (2021).

### v7

@@ 3.2 Transitivity: the Challenge of Identifying Causal Representations
Even with the identifiable $n_{i}$, it is impossible to identify the latent causal variables $z_{i}$ without additional assumptions. To interpret this point, for simplicity, let us only consider the influences of $z_{1}$ and $z_{2}$ on $\mathbf{x}$ in Figure 1. According to the graph structure in Figure 1, assume that $z_{1}:=n_{1}$, $z_{2}:=z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})$ (case 1). We then consider the graph structure shown in the left column of Figure 2, where the edge $z_{1}\rightarrow z_{2}$ has been removed, and assume that $z_{1}:=n_{1}$, $z_{2}:=n_{2}$ and $\mathbf{x}:=\mathbf{f}\circ\mathbf{g}(z_{1},z_{2})$ where $\mathbf{g}(z_{1},z_{2})=[z_{1},z_{1}+z_{2}]$ (case 2). Interestingly, we find that the causal models in case 1 and case 2 generate the same observed data $\mathbf{x}$, which implies that there are two different causal models to interpret the same observed data. Clearly, $z_{2}$ in both the two equivalent causal structures are different, and thus $z_{2}$ is unidentifiable.
@@ 3.3 Scaling Indeterminacy in Causal Representations
The scaling indeterminacy of latent causal variables is also an intrinsic indeterminacy in latent space. Again, for simplicity, we only consider the influence of $z_{1}$ and $z_{2}$ on $\mathbf{x}$ in Figure 1, and assume that $z_{1}:=n_{1}$, $z_{2}:=z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})$. Under this setting, if the value of $z_{1}$ is scaled by $s$, i.e., $s\times z_{1}$, we can easily obtain the same observed data $\mathbf{x}$ by: 1) letting $z_{2}:=\frac{1}{s}\times z_{1}+n_{2}$ and 2) $\mathbf{x}:=\mathbf{f}\circ\mathbf{g}(z_{1},z_{2})$ where $\mathbf{g}(z_{1},z_{2})=[s\times z_{1},z_{1}]$. This indeterminacy is because the scaling of the latent variables $z_{i}$ can be ‘absorbed’ by the nonlinear function from $\mathbf{z}$ to $\mathbf{x}$ and the causal functions among the latent causal variables. Therefore, without additional information to determine the values of the latent causal variables $z_{i}$, it is only possible to identify the latent causal variable up to scaling, not exactly recovering the values. In general, this scaling does not affect identifying the causal structure among the latent causal variables. We will further discuss this point in Section 4.
@@ 3.4 Permutation Indeterminacy in Causal Representations
However, the flexibility and ambiguity above in latent space also offer an advantage, i.e., by pre-defining a causal order, we can avoid constraints like those proposed in Zheng et al. (2018), facilitating the learning of latent causal representations and causal structure. For example, For example, assume that the true latent causal relations for generating data are: size $\rightarrow$ color $\rightarrow$ shape $\rightarrow$ location. During the learning process, we can pre-define a causal order $z_{1},z_{2},z_{3},z_{4}$, without specifying semantic information for the nodes. Since $z_{1}$ is the first node in the predefined causal order, $z_{1}$ is enforced to the semantic information of the first node in the true underlying causal order, i.e., the size. Similarly, $z_{2}$ is enforced to the semantic information of the second node in the true underlying causal order, i.e., the color. Furthermore, the predefined causal order allows us to naturally establish a causal fully-connected graph as depicted in Figure 3, ensuring the estimation of a directed acyclic graph (DAG) in learning causal representation and avoiding DAG constraints, such as those proposed by Zheng et al. (2018). We will further demonstrate how to implement this in the proposed method in Section 6.
@@ 4 Identifiable Causal Representations with Weight-variant Linear Gaussian Models
As discussed in Section 3.2, the key factor that impedes identifiable causal representations is the transitivity in latent space. Note that the transitivity is because the causal influences between the latent causal variables may be ‘absorbed’ by the nonlinear mapping from latent variables $\mathbf{z}$ to the observed variable $\mathbf{x}$. To address this issue, motivated by identifying causal structures with the change of causal influences in observed space (Ghassami et al., 2018; Huang et al., 2019, 2020), we allow causal influences between latent causal variables to be modulated by an additionally observed variable $\mathbf{u}$, represented by the red edge in Figure 4.
Figure 4: The proposed Latent Causal Models. Here we use the ’red’ lines to indicate the changes of weights across $\mathbf{u}$, which is different from the definition of edges (i.e., causal direction) in standard causal model graphs.
Intuitively, variant causal influences between latent causal variables cannot be ‘absorbed’ by an invariant nonlinear mapping from $\mathbf{z}$ to $\mathbf{x}$, breaking the transitivity, resulting in identifiable causal representations. Specifically, we explore latent causal generative models where the observed data $\mathbf{x}$ is generated by the latent causal variables $\mathbf{z}$, allowing for any potential graph structures among $\mathbf{z}$. In addition, there exist latent noise variables $\mathbf{n}$, known as exogenous variables in causal systems, corresponding to latent causal variables. We introduce a surrogate variable $\mathbf{u}$ characterizing the changes in the distribution of $\mathbf{n}$, as well as the causal influences among latent causal variables $\mathbf{z}$. Here $\mathbf{u}$ could be environment, domain, or time index. More specifically, we parameterize the latent causal generative models as follows:
@@
  • In Eq. (2), $\boldsymbol{\lambda_{i}}(\mathbf{u})$ denote the vector corresponding to the causal weights from $\mathbf{z}$, $\mathbf{z}\in\mathbb{R}^{n}$, to $z_{i}$. Specifically, $\boldsymbol{\lambda_{i}}(\mathbf{u})=[\lambda_{1,i}(\mathbf{u}),...,\lambda_{n,i}(\mathbf{u})]$, where each $\lambda_{j,i}$} could be a nonlinear mapping.
@@
In Eq. (2), we assume the latent causal model of each $z_{i}$ satisfies a linear causal model with the causal weights being modulated by $\mathbf{u}$, and we call it weight-variant linear Gaussian models. Therefore, $p({\mathbf{z}|\mathbf{u}})$ satisfies the following multivariate Gaussian distribution:
@@
where $\mathrm{pa}_{i}$ denotes the parent nodes of $z_{i}$. Furthermore, this multi-variate Gaussian can be re-formulated with the following exponential family:
@@
where the parameter $\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u})=[{\Sigma}^{-1}\boldsymbol{\mu};-\frac{1}{2}\operatorname{vec}({{\Sigma}^{-1}})]$, $Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u})$ denotes the normalizing constant, and the sufficient statistic $\mathbf{T}_{\mathbf{z}}(\mathbf{z})=[\mathbf{z};\operatorname{vec}(\mathbf{z}\mathbf{z}^{T})]$ (‘vec’ denotes the vectorization of a matrix). We further denote by $\mathbf{T}_{\text{min}}$ the minimal sufficient statistic and by $k$ its dimension, with $2n\leq k\leq n+(n(n+1))/2$. In particular, $k=2n$ corresponds to the case when $\mathbf{T}_{\text{min}}(\mathbf{z})=[{z_{1}},z_{2},...,z_{n},{z^{2}_{1}},{z^{2}_{2}},...,{z^{2}_{n}}]$, or in other words, there is no edges among the latent variables $\mathbf{z}$, while $k=n+(n(n+1))/2$ corresponds to a full-connected causal graph over $\mathbf{z}$. So with different causal structures over $\mathbf{z}$, the dimension $k$ may vary. For the graph structure in Figure 1, $\mathbf{T}_{\text{min}}(\mathbf{z})=[{z_{1}},z_{2},z_{3},z_{4},{z^{2}_{1}},{z^{2}_{2}},{z^{2}_{3}},{z^{2}_{4}},z_{1}z_{2},z_{1}z_{3},z_{3}z_{4}]$ and $k=11$.
Given the above, the following theorem shows that under certain assumptions on the nonlinear mapping $\mathbf{f}$ and the variability of $\mathbf{u}$, the latent variables $\mathbf{z}$ can be identifiable up to trivial permutation and scaling.
@@
    There exist $2n+1$ distinct points $\mathbf{u}_{\mathbf{n},0},\mathbf{u}_{\mathbf{n},1},...,\mathbf{u}_{\mathbf{n},2n}$ such that the matrix
@@
then the true latent causal variables $\mathbf{z}$, which are learned by matching the true marginal data distribution $p(\mathbf{x}|\mathbf{u})$, are related to the estimated latent causal variables $^{\mathbf{\hat{z}}}$ by the following relationship: $\mathbf{z}=\mathbf{P}\mathbf{\hat{z}}+\mathbf{c},$ where $\mathbf{P}$ denotes the permutation matrix with scaling, $\mathbf{c}$ denotes a constant vector.
@@
The proof can be done according to the following intuition. With the support of assumptions (i)-(iii), one can utilize the identifiability result from nonlinear ICA to identify the latent noise variables $\mathbf{n}$ up to permutation and scaling, i.e., $\mathbf{n}=\mathbf{P}\mathbf{\hat{n}}+\mathbf{c}_{\mathbf{n}}$ where $\mathbf{\hat{n}}$ denotes the recovered latent noise variables obtained by matching the true marginal data distribution, and $\mathbf{c}_{\mathbf{n}}$ notes a constant vector. Next, using a similar technique, we can utilize assumption (iv) to establish a linear transformation between the true latent causal variables $\mathbf{z}$ and the recovered latent causal variables $\mathbf{\hat{z}}$, i.e., $\mathbf{z}=\mathbf{A}(\mathbf{\hat{z}})+\mathbf{c}_{\mathbf{z}}$ where $\mathbf{c}_{\mathbf{z}}$ denotes a constant vector, and $\mathbf{A}$ denotes a liner transformation. Finally, combining the identifiability result for the latent noise variables $\mathbf{n}$ with assumption (v), the linear transformation can be reduced to permutation and scaling, i.e., $\mathbf{z}=\mathbf{P}(\mathbf{\hat{z}})+\mathbf{c}_{\mathbf{z}}$. See Appendix A.1 for details.
@@
Assumptions (i)-(iii) are originally developed by nonlinear ICA (Hyvarinen and Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020; Sorrenson et al., 2020). We here consider unitizes these assumptions considering the following two main reasons. 1) These assumptions have been verified to be practicable in diverse real-world application scenarios Kong et al. (2022); Xie et al. (2022b). 2) Our result eliminates the need to know the exact dimensionality of latent causal or noise variables, which is in contrast to existing methods that require prior knowledge of the dimensionality, due to imposing the two-parameter Gaussian noise variables (Sorrenson et al., 2020). Assumptions (iv)-(v) originally introduced by this work. Intuitively, assumption (iv) is similar to assumption (iii) in the hope that the auxiliary variable $\mathbf{u}$ must have a sufficiently strong and diverse effect on the distributions of latent causal variables $\mathbf{z}$, similar to latent noise variables $\mathbf{n}$.
@@
Among the assumptions above, the only one that cannot be considered weak or natural is assumption (v). Essentially, it is to ensure sufficient changes in causal influence. For simplicity, consider two variables, $z_{1}$ and $z_{2}$, and $z_{2}:=\lambda(\mathbf{u})z_{1}+n_{2}$, where $\lambda(\mathbf{u})=\hat{\lambda}(\mathbf{u})+b$, with $\lambda$ and $\hat{\lambda}$ belonging to the same function class. As a consequence, while the causal influence $\lambda(\mathbf{u})$ changes as a whole across $\mathbf{u}$, there always exists a part $bz_{1}$ that remains unchanged across $\mathbf{u}$. This unchanged term $bz_{1}$ across $\mathbf{u}$ can be absorbed into $\mathbf{f}$ the mapping from $\mathbf{z}$ to $\mathbf{x}$, resulting in a possible solution $\hat{z}_{2}:=\hat{\lambda}(\mathbf{u})z_{1}+n_{2}$, instead of the groundtruth $(\hat{\lambda}(\mathbf{u})+b)z_{1}+n_{2}$, which leads to an unidentifiable outcome. Essentially, this unidentifiable result occurs because a part of a weight remains unchanged across $\mathbf{u}$, a situation we refer to as insufficient change. The objective of assumption (v) is to prevent such insufficient changes. Specifically, by constraining the function class of $\lambda$ such that $\lambda(\mathbf{u}=\mathbf{0})=0$ and thus $\hat{\lambda}(\mathbf{u}=\mathbf{0})=0$, $\lambda(\mathbf{u})$ can not be expressed by $\hat{\lambda}(\mathbf{u})+b$ with $b\neq 0$. Intuitively, assumption (v) suggests that there is a scenario where an existing edge can be removed at a certain $\mathbf{u}$, i.e., $\lambda(\mathbf{u}=\mathbf{0})=0$. By removing the edge, we can ensure that no part of the weight corresponding to the edge remains unchanged. It is important to emphasize that assumption (v) is to limit the function class, so that $\lambda$ do not include a constant term, such as $b$, by assuming $\lambda(\mathbf{u}=\mathbf{0})=0$. Once samples are drawn from this limited function class, the assumption is satisfied, making the observed data generated by these samples effective for inferring latent causal variables. Therefore, we do not need the samples to necessarily include the specific point $\mathbf{u}=\mathbf{0}$, to generate the corresponding observed data, which is then used for inferring the latent causal variables.
Assumption (v) as a Bridge Between Soft and Hard Interventions
Assumption (v) implies that, although a causal edge may exist between two latent variables, there can be a reference condition under which this influence effectively vanishes, i.e., $\lambda(\mathbf{u}=\mathbf{0})=0$. This bears a close connection to hard interventions, which remove causal dependencies by construction. Unlike existing approaches that explicitly rely on hard interventions (Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023), our formulation characterizes a restricted function class for $\lambda(\mathbf{u})$ that enables identifiability without requiring direct access to such interventions. We view this as a conceptual link between soft and hard interventions, while a systematic unification is beyond the scope of this work.
@@
Assumption (v) can arise naturally in biological imaging settings, such as cell imaging experiments where samples are exposed to different small-molecule compounds. In this context, latent variables may represent protein groups, and causal edges correspond to protein–protein interactions (PPIs). Certain compounds are known to disrupt or inhibit specific PPIs, effectively suppressing the corresponding causal influences (Arkin and Wells, 2004; Lu et al., 2020). Such scenarios provide a plausible instance where $\lambda_{i,j}(\mathbf{u}=\mathbf{0})=0$ approximately holds, making the assumption empirically relevant.